In a class of 40 students 15 grew cabbage 10 lettuce 5 both Cabbage and lettuce the rest neither calculate the probability that a randomly selected from class grew cabbage

1
Expert's answer
2022-03-31T06:46:22-0400

Cabbage only = 15 - 5 = 10

Probability a student grew cabbage "=\frac{10}{40}=\frac{1}{4}"